OpenStudy (minion10010164):

Joshua’s math test has 35 questions. If he answered 6/7 of them correctly, how many questions did Joshua answer correctly?

OpenStudy (phi):

can you figure out 6/7 of 35 ?

OpenStudy (minion10010164):

that's a little confusing can you walk me through it?

OpenStudy (phi):

can you figure out 1/2 of 2 ?

OpenStudy (minion10010164):

1??

OpenStudy (phi):

yes. and 1/2 of 4 is 2 notice you get the correct answer if you multiply 1/2 * 2 (or 1/2 * 4) you multiply the fraction times the number. \[ \frac{6}{7} \cdot 35 \]

OpenStudy (phi):

you could do 6 times 35 and then divide by 7 or you could first do 35 divide by 7, then multiply that by 6 (the second way is easier) but both ways give the same answer

OpenStudy (phi):

any luck ?

OpenStudy (minion10010164):

I'm dividing right now

OpenStudy (minion10010164):

Do you know what 210/7 is?

OpenStudy (phi):

yes. but it's easier to first divide 7 into 35 can you do that ?

OpenStudy (minion10010164):

it would be 5 is the answer 30 with nothing left?

OpenStudy (phi):

yes 35/7 is 5 then do 6 times that: 6*5 is 30 notice 210/7 = 30 but multiply 6*35 and then dividing by 7 is harder than 35/7 and then 6* 5 but the answer is 30

OpenStudy (minion10010164):

thank you

OpenStudy (phi):

6/7 of 35 is 30